OpenStudy (mathhelp346):

Using the equation method, 25 is 40% of what number?

OpenStudy (cornzyblack):

Well you want an equation, let x = unknown value So i say (40/100) * x = 25 0.4x = 25 x= 25/0.4 x= 62.5

OpenStudy (mathhelp346):

thanks